A loaf of bread is normally distributed with a mean of22 oz. and a standard deviation of 0.5 oz.
a, what is the probability that a loaf is btwn 22.75and 23.00?
B, 5% of the loafs will weigh less than_____ ounces?
C, assuming 200 loaves are baked on a given day, how many of these will weigh more than 23 ounces? thanks for the help...

Answer by stanbon(75887)

A loaf of bread is normally distributed with a mean of 22 oz. and a standard deviation of 0.5 oz.
a, what is the probability that a loaf is btwn 22.75 and 23.00?
z(22.75) = (22.75-22)/0.5 = 0.75/0.5 = 1.5
z(23) = (23-22)/0.5 = 2
----
P(22.75 < x < 2300) = P(1.5 < z < 2) = 0.0441
--------------------------------------------------------------
B, 5% of the loafs will weigh less than_____ ounces?
Find the z-value with a left tail of 5%:
z = invNorm(0.05) = -1.645
Find the corresponding x value using x = zs + u
x = -1.645*0.5 + 22 = 21.1775
-----------------------------------------------------
C, assuming 200 loaves are baked on a given day, how many of these will weigh more than 23 ounces?
z(23) = 2
P(x > 23) = P(z > 2) = 0.0228
Expected # above 23 = 0.0228*200 = 4.55 loaves
=====================================================
Cheers,
Stan H.
===========